The measure of an angle's supplement is 44 less than the measure of the angle. Find the mesure of the angle and the supplement.
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x+(x-44)=180
2x-44=180
2x=224
x=112
x-44=68

Answer by stanbon(75887)

The measure of an angle's supplement is 44 less than the measure of the angle. Find the mesure of the angle and the supplement.
---------------------------------------------------
Let the measure of "the angle" be "x":
The measure of its supplement is "180-x":
-----
Equation:
180-x = x - 44
2x = 224
x = 112 degrees
-----------
180-x = 68 degrees
---------------
Cheers,
Stan H.